A fuse is made of material that melts when the current density reaches 400 A cm2 . What is the diameter of the fuse wire needed to limit the current to 0.40 A? 


1
Expert's answer
2022-05-13T17:03:28-0400

"\sigma=\frac{4I}{\pi d^2},\implies"

"d=\sqrt\frac{4I}{\pi\sigma}=0.36~mm."